OpenStudy (anonymous):

(-3)^2-2(6^2-9) help

OpenStudy (anonymous):

-45

OpenStudy (anonymous):

thats right it is -45

OpenStudy (anonymous):

but how did u get that>/

OpenStudy (anonymous):

-3^2 = 9 6^2= 36 9-2(36-9) 9-2(27) 9-54= 45

OpenStudy (anonymous):

thank you